LUCIFER Recalibration after Installation on the Telescope

Executive Summary Block
DAYTIME Tests . . . . . . . . . . . . .
  • LUCI1 re-installation on telescope.
  • General function tests
  • MOS tests
  • Field centering and Pupil alignment
  • Check of Grating Tilt (Lambda-V)
  • FlexComp - Verify subset

NIGHTTIME Tests . . . . . . . . . . .
  • Orientation on sky
  • Find Center of Rotation
  • Verification of AGw transform

And if there is time: (NIGHTTIME) . . . . . . . . . . .

  • PM: Photometric ZPs
  • PM: Field illumination
  • PM: Specphot tests
  • Collimation model check
  • Pointing model check
  • Run a science observation

To get Luci1 science ready for the science-block starting April 14 2012

Separated into day-time and night-time activities

  • WS: Walter Seifert
  • JH: Jochen Heidt
  • PB: Peter Buschkamp
  • DT: Dave Thompson
  • LBTO: Unspecified LBTO personnel


On-telescope day-time (closed dome) calibrations, daily schedule:

Monday 09 April 2012

LUCI1 re-installation on telescope. Duration 1.0d
This is NOT the official LUCI transfer procedure !!!

Responsible: LBTO personnel
Comments: Should take a half day but a full day is scheduled.

  • Stop LUCI SW
  • Shut down Sun_V880
  • Move LUCI to telescope
  • Move Sun_V880 to Computer Room B
  • Re-cable fibers for telescope installation
  • Start Sun_V880
  • Start LUCI SW
  • Check communication with instrument
  • Wait for LUCI to return to thermal equilibrium - Detector stable @77K
  • Switch on ROE, wait 2hr

Tuesday 10 April 2012

General function tests Duration: 4h
Responsible: WS
Telescope: @Zenith, Rotator enabled, Lights can be ON
Comments: Use ONLY engineering panels, Main Instrument GUI only AFTER Peter has checked the MOS; start with rotator angle 341deg, repeat for 2-3 other rotator angles

> Basic instrument function tests (from WS email Aug 22, 2011):
  • Stop SW / Re-start SW (optional, restarted in move to telescope)
  • Repeat at ROT ~ 341, 221, 101 degrees:
    • Camera Wheel: Init; move several times between N1.80 & N3.75 camera
    • FW 1 & 2: init, move both wheels front/backwards movements
    • Grating Unit: init, select all gratings/mirror
    • Grating Tilt: set voltages 0V, +-1V
    • Detector focus: move steps to reach neglim & poslim, note # of steps
    • PV: init and move in/out
  • Check that all calib lamps work?

*insert notes here*------------------------------------------

MOS tests Duration: 4h
Responsible: PB
Telescope: @Zenith, Rotator enabled, Lights can be ON

> Basic MOS checkout:
Tests not detailed. Peter's steps to complete the MOS checkout should be documented for future reference.

 #############################################################################
 # Once the MOS tests have been completed (and passed!) and both Walter and  #
 # Peter have given their OKs to continue, you can use the main instrument   #
 # GUI for all remaining tests.  Be sure to start the telescope service in   #
 # the LUCI SW control GUI manually.                                         #
 #############################################################################

Wednesday 11 April 2012

Field centering and Pupil alignment Duration: 5h
Responsible: WS
Telescope: @Zenith, Rotator enabled
LUCI: Imaging, N3.75, K band

> Field centering procedure using FM4
  • Take an image and save it
  • Measure gaps at edges of image
  • FM4: move the mirror in steps to center, notes:
    • Enlarge CompensationMirrorGUI to full screen, use right side of it.
    • M1 -> y, M2-> x; + x steps -> moves image to the right, +y to the top
    • 100 steps corresponds to approximately to 1pix
  • Take images with both cameras, verify centering

> Pupil alignment procedure:
  • Take image to show present alignment of the pupil
  • Adjust FM1 + FM4 (WS: document procedure for future reference)?
  • Repeat until alignment is acceptable

Check Grating Tilts (Lambda-V) Duration: ??
Responsible: JH, DT
Telescope: @Zenith, Rotator enabled, Lights OFF in dome
LUCI: Calib unit in beam, LS0.75_450um slit, grating tilt loop closed.

Spectra: grat*
Throughslit-images: ls*
Done for N1.80-camera only

Test table:
Camera Filter Grating Mask Lamp Voltages to test
N1.80 HKspec 200_H+K LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear z 210_zJHK LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear J 210_zJHK LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear H 210_zJHK LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear K 210_zJHK LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear Ks 150_Ks LS0.75_450um Ar/Xe/Ne 3 different

Flexure Compensation - verification Duration: 3h

Responsible: OKuhn (both JH and DT en route to LBT).
Telescope: EL and Rotator enabled, Lights OFF in dome
LUCI: Calib unit in beam (OK for imaging too, same temp as dome)
Scripts located in /home/observer/LBTO/FlexComp_Verify/
Comment: Verification, only a subset needed
Prereq: Flexure calibrations in place, flexure compensation ON!

> Flexure compensation verification:
  • MANUALLY TURN FLEXURE COMPENSATION ON to verify FC!
  • MCS GUI -> ROTATORS -> MOTION CONTROL, set max speed to 1 deg/sec
  • Set EL, send LUCI rotator to -69 degrees to start
  • Best sequence:
    • one config
    • all ROT positions (script does this, ROT starts/ends @ -69d)
    • 2 EL's: 38 and 72 degrees (mid-points of measurements)
    • next config
  • Process data following ~/LBTO/Flexures/FlexureCheck_Luci1.pdf
  • Copy results files into ~/LBTO/Flexures/FlexComp_Verify
  • MCS GUI -> ROTATORS -> MOTION CONTROL, set max speed to 5 deg/sec
  • Send resulting data to Walter

Camera Filter Grating Mask Lamp Script Name
N3.75 clear K Mirror Spectro Sieve Thermal FC_N375_mirror.cal

Turn on Ar lamp manually!

Camera Filter Grating Mask Lamp Script Name
N1.80 HKspec 200_H+K Spectro Sieve Ar FC_N180_200HK.cal
N1.80 clear J 210_zJHK Spectro Sieve Ar FC_N180_210zJHK.cal

Other notes: subtract a dark or a lamp-off before processing. Should see <~2 pix residuals.
Set Rotator max speed to 1 degree/sec.

Orientation on sky: Duration: 1h
Responsible: DT, JH
Pre-requisite: Orientation done before AGw transform calibration!
Pre-requisite: field with good astrometric coords, good seeing (<1")
Note: Can be done on Persson stds during PM night
Note: Will update further after some MOS acquisitions

> Orientation data
  • Configure for imaging with N3.75, Ks
  • Preset to field (Persson standards, any low-lat field)
  • Take image pair, script: /home/observer/LBTO/SciencePrep/imagepair
  • Repeat for ~3 fields
  • Derive PA offset, update LEFTZEROPOINT (little to no change => DONE!)
  • Update PCSInstrument.conf and RESTART PCS!
  • Repeat observations for ~3 fields
  • Check that orientation is OK

LEFTZEROPOINT updated to -187.814 (from -187.764). Orientation at PAreq=0.00 is -0.029 ± 0.036 degrees (31 measurements).

Note: systematic residuals in the rotator calibration will show as a consistent rotator offset in MOS mask acquisitions that can be "mined" from the event log after science blocks and the rotator ZPs updated. The systematic dX and dY offsets from MOS alignments should also be a measure of the difference between the pointing origin and the mechanical mask center. This should be true for MODS as well as LUCI.

Find Center of Rotation Duration: 1h
Responsible: DT, JH
Pre-requisite: Any dense field
Use pointing mask to also measure mechanical center of masks

> Rotator center
  • Configure for imaging with N3.75, Ks
  • Preset to any dense star field at PA=0
  • Allow to collimate
  • Take image
  • Preset to same field at PA=180
  • Allow to collimate
  • Take image
  • Also do PA=90, 270
  • Repeat on at least two different fields

> Mean coords of matched stars gives rotator center > Mean coords of reference circle gives mechanical center

Verification of AGw transform Duration: 2h
Responsible: ?LBTO (DT, JH to take data)

Once transform is in place, take another set of data to verify. Note: a repeated dice-5 pattern should be well-aligned with rows/columns

And if there is time: (NIGHTTIME) . . . . . . . . . . .

PM: Total system throughput and photometric ZPs: Duration: 7h
Responsible: DT, JH
In all BB-filters
Pre-requisite: photometric conditions

### Standard with a range of colors observed at a range of airmass to
### derive ZP@airmass=1 and extinction, perhaps also color terms.

> Photometric zero points
  • Preset to series of photometric standards (AIRM 1.0…2.0)
  • Take imagepair through each BB filter (K, Ks, H, J, z)
  • Minimum ~5 samples per filter spread in airmass, 10 is better

PM: Field illumination: Duration: 3h
Responsible: DT, JH
N3.75 camera due to new field lens
Pre-requisite: photometric conditions
3h should be enough to do a couple of filters
See /home/observer/LBTO/SciencePrep/Illumination for generic scripts

> Field Illumination
  • Configure for imaging in N3.75 camera
  • Select filter, target (any Persson standard, other fields?)
  • Test DIT, configure NDIT to allow collimation
  • Center source on detector
  • Run at least two scripts per filter:
    • Ddice5_100as_illum.scr (gets the extreme corners, +/-100")
    • D25x_illum1.scr (covers field, monitors extinction)
  • Repeat for all filters (eventually) BB and NB.

Check for a complete science program (SSC in M51, longslit)

Still to add to the plan?
  • Update/check the collimation model (will happen in course of other tests)
  • Update/check the pointing model (will happen in course of other tests)

-- DavidThompson - 27 Feb 2012

-- DavidThompson - 05 Apr 2012
Topic revision: r1 - 05 Apr 2012, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback